# Molecular mechanism of heterogeneous ice nucleation in the atmosphere

### v2

@@ Conclusions
Our MD simulations show that the (110) surface of K-feldspar is responsible for its high ice nucleation efficiency. A comprehensive analysis of alternative feldspar surfaces shows no evidence of enhanced ice nucleation. Our results also provide two key molecular-level insights into this phenomenon. First, we find that the high ice nucleation efficiency of the (110) feldspar surface is explained by its unique ability to structure liquid water into an arrangement closely resembling that on the (110) surface of cubic ice, thereby providing an optimal template for ice nucleation and growth. Second, the ice clusters formed over the (110) feldspar surface are approximately hemispherical, with minor faceting, and mostly composed of cubic ice, a metastable polymorph with respect to hexagonal ice. These findings stand in contrast to previous work, which had suggested that the high-energy feldspar (100) surface provides preferential sites for ice nucleation. Our results contradict this hypothesis and instead support a new interpretation in which atmospheric heterogeneous ice nucleation occurs at feldspar defects exposing patches of the (110) surface. In this new picture, the initial nuclei adopt a cubic ice structure but subsequently transform into hexagonal ice as they grow to macroscopic dimensions. Our work provides a clear molecular-level explanation for the exceptionally high ice nucleation efficiency of K-feldspar, which has been shown to be the dominant ice nucleating particle in the atmosphere.
Acknowledgments
@@
We are grateful for the computational resources provided on the Hyperion cluster at the Donostia International Physics Center (DIPC) and on MareNostrum 5 ACC at the Barcelona Supercomputing Center, granted through the Spanish Supercomputing Network (RES) allocations FI-2024-2-0026 and FI-2024-3-0028. We thank Prof. Hu Qiu from Nanjing University of Aeronautics and Astronautics for valuable suggestions on the data presentation in this work.
Funding:
W.Z. was funded by the ’María de Maeztu’ Units of Excellence program No. CEX2020-001038-M. P.M.P. acknowledges funding from the Marie Sklodowska-Curie Cofund Programme of European Commission project H2020-MSCA-COFUND-2020-101034228-WOLFRAM2.
@@
The scripts used for generating slab surfaces, the dataset for training the MLPs, the input files for MLP training, the final trained MLP, as well as the input files for the MD simulations and DFT calculations reported in this work, are openly available on Zenodo (DOI https://doi.org/10.5281/zenodo.17610332) (?).
Supplementary materials
Methods
Tables S1 to S2
Figs. S1 to S13
References (30-53)

### v3

@@ Conclusion
Our MD simulations show that the (110) surface of K-feldspar is responsible for its high ice nucleation efficiency. A comprehensive analysis of alternative feldspar surfaces shows no evidence of enhanced ice nucleation. Our results also provide two key molecular-level insights into this phenomenon. First, we find that the high ice nucleation efficiency of the (110) feldspar surface is explained by its unique ability to structure liquid water into an arrangement closely resembling that on the (110) surface of cubic ice, thereby providing an optimal template for ice nucleation and growth. Second, the ice clusters formed over the (110) feldspar surface are approximately hemispherical, with minor faceting, and mostly composed of cubic ice, a metastable polymorph with respect to hexagonal ice. These findings stand in contrast to previous work, which had suggested that the high-energy feldspar (100) surface provides preferential sites for ice nucleation. Our results contradict this hypothesis and instead support a new interpretation in which atmospheric heterogeneous ice nucleation occurs at feldspar defects exposing patches of the (110) surface. In this new picture, the initial nuclei adopt a cubic ice structure but subsequently transform into hexagonal ice as they grow to macroscopic dimensions. Although the proposed mechanism relies on the exposure of the (110) plane at defect sites, we note that this surface is common and relatively easy to cleave, ranking just below the (001) and (010) surfaces in terms of cleavage perfection[38]. Our work provides a clear molecular-level explanation for the exceptionally high ice nucleation efficiency of K-feldspar, which has been shown to be one of the key ice nucleating particles in the atmosphere.
Methods
Generating feldspar surface terminations
All possible terminations of the (001), (010), (100), (110), (10), and (01) planes were generated using the SlabGenerator module of the pymatgen package[49]. A cutoff of 1 Å along the surface normal direction was applied to determine whether atoms lie within the same atomic plane, which was used to filter the generated terminations. There is extensive experimental evidence for the absence of reconstructions in feldspar surfaces and thus we only consider flat unreconstructed surfaces[8, 5, 6]. Hydrogen atoms were inserted near undercoordinated oxygen atoms at the surfaces, in accordance with recent atomic force microscopy observations, which show full surface hydroxylation[8]. In addition, surface terminations that did not preserve the integrity of SiO4 and AlO4 polyhedra or failed to maintain charge neutrality were discarded. Consequently, a total of 13 distinct surface terminations were obtained, including 2 terminations for the (001) plane, 2 for the (010) plane, 2 for the (100) plane, 3 for the (110) plane, 3 for the (10) plane, and 1 for the (01) plane (see Table S1 and S2). Among these, all terminations of the (001), (010), and (100) planes correspond to those previously reported, namely, the (001)-, (001)-, (010)-, (010)-, (100)-, and (100)- surfaces[39, 33]. Three terminations were identified for the (110) plane, each of them involving the breaking of eight bonds per unit cell (Table S1). In one of the terminations, three oxygen atoms of a SiO4 tetrahedron are disconnected from their neighboring two Si and one Al atoms, leaving only one oxygen atom bonded to the surface. This configuration is likely unstable and was thus denoted as the (110)- termination. In another case, the cleavage requires breaking three bonds of a SiO4 tetrahedron, which is energetically more unfavorable than cases where only one or two bonds are broken. This termination was therefore labeled as (110)-, while the remaining, more stable termination was designated as (110)-. The (10) plane is a mirror-related counterpart of the (110) plane, and consequently exhibits atomic-scale structural features similar to those of the (110) surface. Therefore, the terminations of the (10) plane were named following the same convention, namely, (10)-, (10)-, and (10)-. The (01) plane has only a single termination, which is therefore designated as (01)-.
The large K-feldspar slabs used for molecular dynamics (MD) simulations expose two identical surfaces along the -axis, with atomic-scale structures of both surfaces corresponding to one of the terminations listed above. One of the surfaces is in direct contact with water, forming the water–feldspar interface. A vacuum layer is introduced along the -axis between the water layer and the opposite surface of the feldspar slab to eliminate spurious interactions.
Electronic structure calculations
Plane-wave DFT calculations were performed using the Quantum ESPRESSO suite for electronic structure calculations v6.4.1[10, 9]. We used the Strongly Constrained and Appropriately Normed (SCAN) exchange and correlation functional[46, 45] as implemented in the LIBXC 4.3.4 library[28]. We employed norm-conserving, scalar-relativistic pseudopotentials[12] for K, Al, Si, O, H parametrized using the PBE[32] functional with 9, 11, 4, 6, and 1 valence electrons, respectively. Kinetic energy cutoffs of 110 and 440 Ry were used for the wave functions and the charge density. We used only the -point in our calculations. All other parameters were set to their default values in Quantum ESPRESSO[10, 9].
Machine learning interatomic potential
The smooth-edition of the Deep Potential methodology developed by Zhang et al., as implemented in DeePMD-kit v2.10.0, was used to train the machine-learning interatomic potentials (MLPs)[59, 55, 60]. An active learning strategy was employed during the training process, as shown in Fig. S12. First, MD simulations driven by our previous MLP (Ref. Piaggi et al. [33]) were performed to generate a series of configurations of the water-feldspar interfaces, covering all 13 K-feldspar terminations. The energies and forces for these configurations were then calculated using SCAN DFT to expand our dataset reported in Ref. Piaggi et al. [33]. The resulting dataset, which included new configurations and their corresponding energies and atomic forces, was used to train a set of four MLPs. Based on the newly trained MLPs, additional configurations were explored, and this cycle was repeated iteratively until a high-accuracy MLP with SCAN-level precision was obtained. Figures S13 and S14 show that our MLPs have excellent and uniform accuracy across all studied terminations. Settings of DeePMD-kit are identical to those used in Ref. Piaggi et al. [33].
MD simulations
We performed MD simulations and enhanced sampling simulations using LAMMPS[35, 48] interfaced with the PLUMED plugin[50, 47]. All simulations employed the custom-trained MLPs described above to model atomic interactions. We used the standard atomic weights for the masses of all elements (K, Al, Si, O) except H, for which we used a mass of 2 grams/mol, in order to improve the stability of the integration of the equations of motion. Simulation boxes with periodic boundary conditions in all directions were used. The time step for the integration of the equations of motion was 0.5 fs. The simulation models were first energy-minimized using the conjugate gradient (CG) method and then equilibrated using MD simulations in the ensemble. A stochastic velocity-rescaling thermostat[2] with a 0.1 ps relaxation time was used to control the temperature. For studies of water structure on multiple surfaces, simulations were performed at 290 K (supercooling =18 K) for more than 3 ns. During the simulations, the K, Al, and Si atoms were constrained to their initial positions by a harmonic spring with force constant of 20 . The trajectory was saved every 1000 steps.
As standard MD simulations cannot capture the nucleation process within affordable simulation times, we employed an enhanced sampling method, namely steered MD simulations guided by the Steinhardt order parameter[43, 51], as implemented in PLUMED[50, 47]. The oxygen atoms of water molecules were used to calculate the Steinhardt order parameter, which quantifies local structural ordering. A moving restraint with a force constant of was applied to this collective variable. The simulations were performed at 300 K (supercooling =8 K). The target value of the restraint was gradually increased from 0.05 to 0.2 over a 5 ns simulation. The order parameter was recorded every 1000 steps to monitor the progress of nucleation throughout the simulation, and the trajectory was saved every 10000 steps. After the steered MD simulations, configurations were extracted from the trajectory and used as the initial models for seeding simulations without bias. A very similar steered MD protocol was extensively validated for studying ice-binding surfaces in Ref. Naullage et al. [30]. The simulations reported in that work show that while using biased simulations may result in a different polymorph than in the unbiased simulations, this impacts mostly the ice that forms far from the surface.
The free energy profile of ice nucleation was calculated using the umbrella sampling method. The Steinhardt parameter was used as collective variable, and 41 windows in the range with a width of 0.0015 were used. In each window, the parameter was restrained by a harmonic potential with a spring constant of . Each simulation was performed for 3 ns. The time series of the parameter for all windows were then used to construct the free energy profile via the weighted histogram analysis method (WHAM), following the implementation of Grossfield[26, 37, 11].
Visualization and analysis of the MD trajectories were performed using VMD[18]. To identify ice-like atomic environments we employed the Identify Diamond Structure modifier in OVITO[44] using the oxygen atoms. We quantified the number of ice-like water molecules by counting the number of oxygen atoms classified as having a diamond-like structure, and we ignored their first and second neighbors with imperfect structures, which are sometimes included in this type of calculation. The results are fully consistent with the values obtained using the Polyhedral Template Matching modifier[27, 34].
Obtaining thermodynamic properties using classical nucleation theory
The effective surface free energy for the heterogeneous nucleation process can be computed from an equation that closely follows its counterpart for homogeneous nucleation, namely,
where is an effective surface free energy that combines information about the water-ice surface free energy and the contact angle, is the density of ice, and is the chemical potential difference between liquid water and ice. and are easily computed and were taken from previous SCAN-MLP-based simulations reported in ref. Piaggi et al. [34]. According to classical nucleation theory, is related to the nucleation free energy barrier, , as follows,
Another way to characterize the ice nucleation process is through the heterogeneous nucleation rate, which can be calculated as,
where molecules/nm2 is the density of liquid water at the first layer in contact with the substrate’s surface, is the Zeldovich factor, is the attachment rate of molecules to the critical nucleus, with being the Boltzmann constant and the temperature. and were taken from ref. Piaggi et al. [34].
Estimation of nucleation rates in experiments
We estimated the experimental ice nucleation rate using Fig. 2A of Ref. Kiselev et al. [23]. In this figure, the nucleation sites of individual ice crystals were plotted during eight subsequent nucleation–evaporation cycles on the defects of a feldspar (001) cleavage surface within 100 s. Approximately 200 sites were marked over the eight cycles in the observed region. The area of the observed (001) cleavage surface is . Clearly, the exposed (110) surface area is much smaller than that of the (001) surface, and thus the area of (110) surface was estimated using a scaling factor, , such that the effective exposed area of (110) surface is . Therefore, the ice nucleation rate, , can be estimated as
where represents the average number of nucleation sites per cycle, and is the duration of each nucleation–evaporation cycle. The estimated nucleation rate is . The scaling factor is much smaller than 1, and assuming to be in the order of to , we computed the range , which is highlighted in Fig. 3I.
Acknowledgements
@@
We are grateful for the computational resources provided on the Hyperion cluster at the Donostia International Physics Center (DIPC) and on MareNostrum 5 ACC at the Barcelona Supercomputing Center, granted through the Spanish Supercomputing Network (RES) allocations FI-2024-2-0026 and FI-2024-3-0028. We thank Prof. Hu Qiu from Nanjing University of Aeronautics and Astronautics for valuable suggestions on the data presentation in this work. W.Z. was funded by the ’María de Maeztu’ Units of Excellence program No. CEX2020-001038-M. P.M.P. acknowledges funding from the Marie Sklodowska-Curie Cofund Programme of European Commission project H2020-MSCA-COFUND-2020-101034228-WOLFRAM2.
@@
The scripts used for generating slab surfaces, the dataset for training the MLPs, the input files for MLP training, the final trained MLP, as well as the input files for the MD simulations and DFT calculations reported in this work, are openly available on Zenodo (DOI https://doi.org/10.5281/zenodo.17610332)[61].